Use the method of substitution to solve the system of linear equations.
3a=7-b
3a=2-b
__A. There is one solution. The solution is __.
__B. There are infinitely many solutions.
__C. there are no solutions.

Answer by jim_thompson5910(35256)

Hint:
3a = 7 - b
3a = 2 - b

basically says

some number = 7 - b
same number = 2 - b